OpenStudy (anonymous):

can someone please help what are the complex zeros of the polynomial x^2+4x-32

OpenStudy (tkhunny):

Have you considered the Quadratic Formula?

OpenStudy (tkhunny):

Have you considered factoring?

OpenStudy (muzzack):

Start out with simplifying it 1 Multiply Terms \(\color{lime}{x^2+4x−32}\) 2 Simplify Exponents \(\color{lime}{x^2+4x−32}\)

OpenStudy (anonymous):

no im not sure how to do that can u show me plesee

OpenStudy (muzzack):

@tkhunny could you explain to him

OpenStudy (anonymous):

I know how to find the zeros...I am not sure what "complex zeros" are though....

OpenStudy (accidentalaichan):

This is funny because I just finished this module in FLVS. we can factor this problem into (x^2 - 8)(x^2 + 4) and then further into (x - √8)(x + √8)(x^2 + 4). Right away, we can tell 2 of the roots are ±√8 or ±2√2. You may already know the roots of x^2 + 4, but if you don't you can find them with the quadratic formula: (-b ± √(b^2 - 4ac)) / (2a) (±√(-4*1*4)) / (2*1) (±i√16) / 2 ±4i / 2 ±2i Thus, we have found all 4 roots. The zeros of the function are x = ±2√2, ±2i.

OpenStudy (anonymous):

@AccidentalAiChan u did thanks so much you know what was on the DBA for module 4 u almost done with module 5

OpenStudy (accidentalaichan):

Actually, I'm in Module 10. They will just ask you how to use the quadratic formula and the graphing terms.

OpenStudy (tkhunny):

All zeros are Complex. Just find ALL of them. In this case, there are none that are not Real. x^2 + 4x - 32 = (x+8)(x-4) and we see x = -8 or x = 4. If you continue to search for a solution with an "i" in it, you will search for a very long time. There isn't one. The Quadratic Formula would have found these, as well.

OpenStudy (anonymous):

@mathmale can u help

OpenStudy (anonymous):

@ParthKohli u know how

OpenStudy (anonymous):

i dont understand